Find the maximum and minimum of e-x2–v? (x² + 2y) on the disk x2 + y2 < 2.
Answer #1

* Given that let, f(x,y) = e-22y? (x²+242) fa = e-zzy? (2n) +822yh), e-x2y? .(-22) -2x[1-23 242] fy (44) + (27262).enazy? .(-The and Critical point bei (0,0) from x?ty2 s2 be! co, &2) and (I/2, o } eo at at 2 0,0), f(x,y) -(x,y) se 1 C0,1V2), fcx,y)=